A ball is thrown upward at 16m/s. How long does it take the ball to reach its maximum height?
1
Expert's answer
2020-05-05T18:40:43-0400

The time when a ball reaches a top

"t=\frac{v_0}{g}=\frac{16\:\rm m/s}{9.8 \:\rm m/s^2}=1.63\:\rm s"